# SparseDM: Toward Sparse Efficient Diffusion Models

### v2

@@ Abstract
Diffusion models have been extensively used in data generation tasks and are recognized as one of the best generative models. However, their time-consuming deployment, long inference time, and requirements on large memory limit their application. In this paper, we propose a method based on the improved Straight-Through Estimator to improve the deployment efficiency of diffusion models. Specifically, we add sparse masks to the Convolution and Linear layers in a pre-trained diffusion model, then use design progressive sparse model training in the fine-tuning stage, and switch the inference mask on, which supports sparse inference. Experiments on a Transformer-based and UNet-based diffusion model demonstrate that our method reduces MACs by $50\%$ while increasing FID by only 1.22 on average. Sparse models are accelerated by approximately 1.2x on the GPU. Under other MACs conditions, the FID is also lower than 1$\sim$137 compared to other methods.
@@ 1 Introduction
Diffusion Models (DM) Song et al. (2021); Karras et al. (2022) have been one of the core generative modules in various computer vision tasks with their advantages on both sample quality and mode coverage. Generally, they are composed of two stages : a forward/diffusion process that perturbs the data distribution to learn the time-dependent score functions, and a reverse/sampling process that generates data samples from a prior distribution in an iterative manner. However, their slow inference speed and heavy computational load during the inference process inevitably restrict their applications on most mobile devices.
To reduce the computational load in the inference process of diffusion models, various methods have been proposed to minimize the number of inference steps, such as the training-free samplers Bao et al. (2022); Zhao et al. (2023); Lu et al. (2022, 2023); Zheng et al. (2023); Chen et al. (2023) and the distillation methods Salimans and Ho (2022); Luo et al. (2023a, b). However, their sample quality is still unsatisfactory as a few sampling steps cannot faithfully reconstruct the high-dimensional data space, e.g., the image or video samples. This problem is even more pronounced on mobile devices with limited computing capabilities. Simultaneously, a few works explore reducing the Multiple-Accumulate operations (MACs) at each inference step Bolya and Hoffman (2023); Fang et al. (2023a). However, their work cannot be accelerated on GPUs. Since DM is an intensive model parameter calculation, NVIDIA Ampere architecture GPU supports a 2:4 sparse (4-weights contain 2 non-zero values) model calculation, which can achieve nearly 2 times calculation acceleration Pool and Yu (2021); Mishra et al. (2021). Although structural pruning Fang et al. (2023a) has been used in DM, 2:4 structured sparsity inference has not been implemented. Our method aims to reduce MACs at each step through 2:4 and other scale sparsity.
@@
We use a state-of-the-art Transformer-based DM, named U-ViT Bao et al. (2023), to analyze the shortcomings of existing 2:4 structured sparse tools. Also, we manually design a half-size DM network to test FID and MACs. From Table 1, we can see that simply reducing the U-ViT model parameters by nearly $50\%$ may cause FID to collapse catastrophically. Automatic Sparsity (ASP) Pool and Yu (2021); Mishra et al. (2021)) for model sparse training consumes a lot of GPU time, but FID is poor. Therefore, for DM, we need to redesign the pruning method to reduce the amount of calculation and maintain FID as much as possible.
Table 1: The results of manually redesigning the UViT model and sparse pruning the UViT model using ASP. UViT Small is from the U-ViT method. Half UViT Small is set to half the depth and number of heads of UViT Small. ASP is the method of sparse pruning from Nvidia.
Datasets Models Methods FID MACs (G) CIFAR10 32x32 UViT Small U-ViT 3.20 11.34 Half UViT Small U-ViT 678.20 5.83 UViT Small ASP 319.87 5.76 CelebA 64x64 UViT Small U-ViT 2.87 11.34 Half UViT Small U-ViT 441.37 5.83 UViT Small ASP 438.31 5.76
This paper proposes a new method for implementing 2:4 structured sparsity and other scale sparse inference. (1) Sparse masks are applied to the Convolution and Linear layers. (2) The STE method improved by sparse regularization is used to train the sparse model. (3) Progressive sparse models are trained by transferring knowledge from a dense model to a sparse model.
@@
  • We propose a progressive sparse finetuning method that achieves 2:4 structured sparse inference and other scale sparse pruning for DMs with a Transformer or UNet backbones.
  • We conduct experiments on four datasets. The average FID of 2:4 sparse DM is only increased by 1.22 compared with the dense model. The inference acceleration on the GPU is approximately 1.2x.
  • Testing nine scales, with similar MACs for each scale, our sparse inference FID is $1\sim 137$ lower than other methods.
@@ 2.1 Diffusion Models
where ${\bm{x}}_{t}$ is input data at $t$. The backward process is formalized as:
@@
where $\alpha_{t}$ and $\beta_{t}$ are the noise schedule at $t$, $\alpha_{t}+\beta_{t}=1$. $\overline{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$, and ${\bm{\epsilon}}$ is the standard Gaussian noises added to ${\bm{x}}_{t}$. The DM training is a task of minimizing the noisy prediction errors, expressed as:
$\min_{\mathcal{W}}\mathcal{L}(\mathcal{W};\mathcal{D}),$ (3)
where $\mathcal{D}$ is dataset, $\mathcal{L}$ is loss function, $\mathcal{W}$ is dense weights. The DM inference, data generation, is expressed as:
${\bm{y}}_{t}=\mathcal{F}(\mathcal{W};{\bm{x}}_{t}).$ (4)
where $\mathcal{F}$ is trained DM, ${\bm{y}}_{t}$ is the output of DM inference.
@@ 2.2 Sparse Pruning
The one-shot training method is easy to use, and the steps are as follows Mishra et al. (2021); Pool and Yu (2021): Train a regular dense model. Prune the weights on the fully connected and convolutional layers in a 2:4 sparse mode. Retrain the pruned model. The one-shot pruning is expressed as:
$\begin{split}\widetilde{\mathcal{W}}&=\mathcal{W}\odot\mathcal{M},\end{split}$ (5)
@@
One-shot sparse pruning can cover most tasks and achieve speedup without losing accuracy. However, for some difficult tasks that are sensitive to changes in weight values, doing sparse training for all weights at once will result in a large amount of information loss Han et al. (2015). With the same number of tuning iterations, progressive sparse training can achieve higher model accuracy than one-shot sparse training. Suppose there are $k$ masks $[\mathcal{M}_{1},\mathcal{M}_{2},...,\mathcal{M}_{k}]$, the progressive sparse process is formalized as:
@@
The reason why traditional progressive sparse training methods work well is to reuse the knowledge of dense models as much as possible. However, this progressive sparse training method has only proven effective when the training data distribution is stable, such as when training classification models.
@@ 2.3 Distribution Shift on DM Training and Sparse Pruning
Robust Fairness Regularization (RFR) Jiang et al. (2024) improved that distribution shift can be transformed as data perturbation, and data perturbation and model weight perturbation are equivalent for classifier models. Assuming that the perturbation of the label is not considered, the equivalent of data perturbation and model weight perturbation is formalized as:
@@
The training process of diffusion models generally involves only the distribution shift of the noisy data. Existing sparse pruning methods generally do not consider the distribution shift of noisy data but only the model’s weight changes. Inspired by RFR’s conclusions, we convert the distribution shifts caused by changes in model weights into distribution shifts caused by data changes for the diffusion model’s training process.
@@ 3 Sparse Finetuning DM
Figure 1: Framework Overview. Training is the process of training with a progress sparse mask.
@@ 3.1 Straight-Through Estimator for Sparse Training
A straightforward solution for training an $N$:$M$ sparsity network is to extend the Straight-through Estimator (STE) Bengio et al. (2013) to perform online magnitude-based pruning and sparse parameter updating. The concrete formula is described as:
@@
where $\gamma$ is learning rate, $g$ is gradient. In STE, a dense network is maintained during the training process. In the forward pass, the dense weights $\mathcal{W}$ are projected into sparse weights $\widetilde{\mathcal{W}^{(N:M)}}=S(\mathcal{W},N,M)$ satisfying $N$:$M$ sparsity. $S(\cdot)$ is projection function. The sparse DM training task is expressed as:
@@
To mitigate the negative impact of the approximate gradient calculated by vanilla STE, we incorporate sparse mask information into the backward propagation process. The updated formula is modified as follows:
@@
where $\lambda$ is coefficient. To obtain extremely sparse DM, jumping directly from the dense model to the extremely sparse model, such as 1:32 ($\frac{1}{32}=0.03125$) sparsity, for training and using the extremely sparse weight gradient for reverse propagation can easily lead to the knowledge loss of dense model or even collapse of the trained model.
3.2 Progressive Sparse Finetuning Diffusion Model
Traditional progressive sparse training works well for extremely sparse model training. However, its training data distribution does not change. Based on the analysis of two distribution shifts in subsection 3.3, only one distribution shift is generally optimized during training. If two distribution shifts are optimized at the same time, such as switching the sparsity rate when training a DM, the optimization will fail. If a highly sparse DM is trained from a dense DM, the traditional progressive sparse training method needs to be changed. The sparsity rate can only be fixed during the DM training. Similar to the traditional progressive sparse training method, we transfer knowledge to the sparse model through the samples generated by the dense model to accelerate the sparse training of DM and reuse the knowledge of the dense model as much as possible.
The prediction of a dense model transfers the original knowledge to a sparsely structured model and does not involve sampling compression; that is, the number of sampling steps is consistent by default. During sparse training, a neural network $\mathcal{F}_{\mathcal{W}}({\bm{x}}_{t},t)$ is trained to predict the noise in ${\bm{x}}_{t}$ by minimizing the L2 loss between them. The loss is formulated as follows:
@@ 3.2 Progressive Sparse Finetuning Diffusion Model
The default pre-trained model is a dense model. The overall loss function of progressive sparse training is composed of the original dense task loss, a diffusion loss that optimizes the diffusion model. It can be formulated as:
$\mathcal{L}_{\mathrm{sparse}}^{i}=\mathcal{L}_{\mathrm{dense}}+\lambda\mathcal{L}_{\mathrm{diff}},~{}~{}i\in[1,k]$ (13)
where $\lambda$ is the loss weight to balance the losses. We simply set $\lambda=1$ in all experiments. $k$ is the sparse mask number.
@@
$\begin{split}\mathcal{W}_{t+1}&\leftarrow\mathcal{W}_{t}-\gamma_{t}g(\widetilde{\mathcal{W}^{(N_{i}:M_{i})}_{t}}),~{}~{}i\in[1,k]\end{split}$ (15)
@@
$\begin{split}\mathcal{W}_{t+1}&\leftarrow\mathcal{W}_{t}-\gamma_{t}\left(g(\widetilde{\mathcal{W}^{(N_{i}:M_{i})}_{t}})+\lambda_{W}(\mathcal{W}_{t}-\widetilde{\mathcal{W}^{(N_{i}:M_{i})}_{t}}))\right),~{}~{}i\in[1,k]\end{split}$ (16)
@@
The structured sparse function provides fully connected layers and convolutional layers with 2:4 sparse weights to achieve 2:4 sparse acceleration on the GPU. If their weights are pruned ahead of time, these layers can be accelerated using structured sparse functions on the GPU. In order to test real inference acceleration on an NVIDIA GPU, we used the 2:4 sparse operator in the acceleration library provided by NVIDIA. NVIDIA Ampere architecture GPUs have implemented CUDA operators to accelerate the multiplication of such matrices. 2:4 sparsity inference is expressed as:
@@
Transposable masks Hubara et al. (2021) suggest that a weight matrix and its transpose can be simply pruned by multiplying binary masks, so the backward pass shares the same sparse weight matrix with the forward pass. Transposable masks are one of the important technologies for NVIDIA Ampere architecture GPUs to accelerate sparse matrices.
@@ 4.1 Experimental Settings
Baseline Methods. We compare the performance of our method with the following baseline algorithm at 2:4 or other sparsity. Diff-Pruning Fang et al. (2023a) is a method of structural pruning of diffusion models. ASP (Automatic SParsity) Pool and Yu (2021); Mishra et al. (2021) is a tool that enables 2:4 sparse training and inference for PyTorch models provided by Nvidia. Nvidia developed a simple training workflow that can easily generate a 2:4 structured sparse network that matches the accuracy of the dense network. To compare the performance of other sparse masks, we give the results of inference with different sparsity rates. Finally, we provide the ablation study results of STE-based for the training process and traditional progressive sparse training. We give the ablation study results of untrained and STE-trained masks for the inference process.

### v3

@@ Abstract
Diffusion models represent a powerful family of generative models widely used for image and video generation. However, the time-consuming deployment, long inference time, and requirements on large memory hinder their applications on resource constrained devices. In this paper, we propose a method based on the improved Straight-Through Estimator to improve the deployment efficiency of diffusion models. Specifically, we add sparse masks to the Convolution and Linear layers in a pre-trained diffusion model, then transfer learn the sparse model during the fine-tuning stage and turn on the sparse masks during inference. Experimental results on a Transformer and UNet-based diffusion models demonstrate that our method reduces MACs by $50\%$ while increasing FID by only 0.44 on average. Sparse models are accelerated by approximately 1.2x on the GPU. Under other MACs conditions, the FID is also lower than 1 compared to other methods.
@@ 1 Introduction
Diffusion Models (DM) (Song et al., 2021; Karras et al., 2022) have been one of the core generative modules in various computer vision tasks. Generally, they are composed of two parts: a forward/diffusion process that perturbs the data distribution to learn the time-dependent score functions, and a reverse/sampling process that generates data samples from a prior distribution in an iterative manner. Though diffusion models have advantages on both sample quality and mode coverage over other competitors, their slow inference speed and heavy computational load during the inference process inevitably restrict their applications on most mobile devices.
To reduce the computational load in the inference process of diffusion models, various methods have been proposed to minimize the number of inference steps, such as the training-free samplers (Bao et al., 2022; Zhao et al., 2023; Lu et al., 2022; 2023; Zheng et al., 2023) and the distillation methods (Salimans & Ho, 2022; Luo et al., 2023b; a). However, their sample quality is still unsatisfactory as a few sampling steps cannot faithfully reconstruct the high-dimensional data space, e.g., the image or video samples. This problem is even more pronounced on mobile devices with limited computing capabilities. Simultaneously, a few works explore reducing the Multiple-Accumulate operations (MACs) at each inference step (Bolya & Hoffman, 2023; Fang et al., 2023b). However, their work cannot be accelerated on GPUs. Since DM is an intensive model parameter calculation, NVIDIA Ampere architecture GPU supports a 2:4 sparse (4-weights contain 2 non-zero values) model calculation, which can achieve nearly 2 times calculation acceleration (Pool & Yu, 2021; Mishra et al., 2021). Although structural pruning (Fang et al., 2023b) has been used in DM, 2:4 structured sparsity inference has not been implemented. Our method aims to reduce MACs at each step through 2:4 and other scale sparsity.
@@
We use a state-of-the-art Transformer-based DM, named U-ViT (Bao et al., 2023), to analyze the shortcomings of existing 2:4 structured sparse tools. Also, we manually design a half-size DM network to test FID and MACs. From Table 1, we can see that simply reducing the U-ViT model parameters by nearly $50\%$ may cause FID to collapse catastrophically. Automatic Sparsity (ASP) (Pool & Yu, 2021; Mishra et al., 2021)) for model sparse training consumes a lot of GPU time, but FID is poor. Therefore, for DM, we need to redesign the pruning method to reduce the amount of calculation and maintain FID as much as possible.
Table 1: The results of manually redesigning the U-ViT model and sparse pruning the U-ViT model using ASP. UViT Small is from the U-ViT method. Half UViT Small is set to half the depth and number of heads of UViT Small. ASP is the method of sparse pruning from Nvidia.
Datasets Models Methods FID MACs (G)
CIFAR10 32x32 UViT Small U-ViT 3.97 11.34
Half UViT Small U-ViT 678.20 5.83
UViT Small ASP 319.87 5.76
CelebA 64x64 UViT Small U-ViT 3.30 11.34
Half UViT Small U-ViT 441.37 5.83
UViT Small ASP 438.31 5.76
This paper proposes a new method for implementing 2:4 structured sparsity and other scale sparse inference. (1) Masks with different sparsity rates are applied to convolutional and linear layers. (2) Sparse regularization is added to back-propagation, improving the STE method to train sparse models. (3) Knowledge is gradually transferred from dense models to sparse models to improve models’ performance with high sparsity rates.
@@
  • We propose a transfer learn sparse masks method that achieves 2:4 structured sparse inference and other scale sparse pruning for DMs with a Transformer or UNet backbones.
  • We conduct experiments on four datasets. The average FID of 2:4 sparse DM is only increased by 0.44 compared with the dense model. The inference acceleration on the GPU is approximately 1.2x.
  • Testing nine scales, with similar MACs for each scale, our sparse inference FID is $1$ lower than other methods.
@@ 2.1 Diffusion Models
where ${\bm{x}}_{t}$ is input data at $t$. In the backward process, the mean network of the denoising transition probability is formalized as:
@@
where $\alpha_{t}$ and $\beta_{t}$ are the noise schedule at $t$, $\alpha_{t}+\beta_{t}=1$. $\overline{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$, and ${\bm{\epsilon}}$ is the standard Gaussian noises added to ${\bm{x}}_{t}$. The DM training is a task of minimizing the noisy prediction errors, expressed as $\min_{\mathcal{W}}\mathcal{L}(\mathcal{W};\mathcal{D})$, where $\mathcal{D}$ is dataset, $\mathcal{L}$ is loss function, $\mathcal{W}$ is dense weights. The DM inference, data generation, is expressed as ${\bm{y}}_{t}=\mathcal{F}(\mathcal{W};{\bm{x}}_{t})$, where $\mathcal{F}$ is trained DM, ${\bm{y}}_{t}$ is the output of DM inference.
@@ 2.2 Sparse Pruning
The one-shot training method is easy to use, and the steps are as follows (Mishra et al., 2021; Pool & Yu, 2021): Train a regular dense model. Prune the weights on the fully connected and convolutional layers in a 2:4 sparse mode. Retrain the pruned model. The one-shot pruning is expressed as:
$\widetilde{\mathcal{W}}=\mathcal{W}\odot\mathcal{M},$ (3)
@@
One-shot sparse pruning can cover most tasks and achieve speedup without losing accuracy. However, for some challenging tasks that are sensitive to changes in weight values, doing sparse training for all weights at once will result in a large amount of information loss (Han et al., 2015). With the same number of tuning iterations, progressive sparse training can achieve higher model accuracy than one-shot sparse training. Suppose there are $k$ masks $[\mathcal{M}_{1},\mathcal{M}_{2},...,\mathcal{M}_{k}]$, the progressive sparse process is formalized as:
@@
The reason why traditional progressive sparse training methods work well is to reuse the knowledge of dense models as much as possible. However, this progressive sparse training method has only proven effective when the training data distribution is stable, such as when training a CNN classification model.
@@ 2.3 Distribution Shift on DM Training and Sparse Pruning
Robust Fairness Regularization (RFR) (Jiang et al., 2024) improved that distribution shift can be transformed as data perturbation, and data perturbation and model weight perturbation are equivalent for classifier models. Assuming that the perturbation of the label is not considered, the equivalent of data perturbation and model weight perturbation is formalized as:
@@
The training process of DMs generally involves only the distribution shift of the noisy data. Existing sparse pruning methods generally do not consider the distribution shift of noisy data but only the model’s weight changes. Inspired by RFR’s conclusions, we convert the distribution shifts caused by changes in model weights into distribution shifts caused by data changes for the DM’s training process. The DM’s sparse ratio is fixed, and then the knowledge of the dense model is transferred to the sparse model.
@@ 3 Sparse Finetuning DM
Figure 1: Framework Overview. This includes the process of transfer learning sparse models.
@@ 3.1 Straight-Through Estimator for Sparse Training
A straightforward solution for training an $N$:$M$ sparsity network is to extend the Straight-through Estimator (STE) (Bengio et al., 2013) to perform online magnitude-based pruning and sparse parameter updating. The concrete formula is described as:
@@
where $\gamma$ is learning rate, and $g$ is the gradient. In STE, a dense network is maintained during the training process. In the forward pass, the dense weights $\mathcal{W}$ are projected into sparse weights $\widetilde{\mathcal{W}^{(N:M)}}=S(\mathcal{W},N,M)$ satisfying $N$:$M$ sparsity, and here $S(\cdot)$ is a projection function. The sparse DM training task is expressed as:
@@
We incorporate sparse mask information into the backward propagation process to mitigate the negative impact of the approximate gradient calculated by vanilla STE. The updated formula is modified as follows:
@@
where $\lambda_{W}$ is a tunable hyperparameter.
Traditional progressive sparse training works well for highly sparse model training (Han et al., 2015; Pool & Yu, 2021). However, its training data distribution does not change. According to the analysis of the two distribution shifts in Section 2.3, only one distribution shift is usually optimized when training using stochastic gradient descent or its improved algorithms. The optimization will fail if two distribution shifts are optimized simultaneously, such as switching the sparsity rate when training a DM. Directly training an extremely sparse model, such as 1:32 ($\frac{1}{32}=0.03125$) sparsity, initialized by dense model weight and using extremely sparse weight gradients for reverse propagation may cause the sparse model to collapse. In this process, knowledge of the dense model is easily lost.
3.2 Transfer Learn Sparse Diffusion Models
To train a highly sparse DM from a dense DM, we fix the sparsity rate during DM training and transfer the knowledge to the sparse model via samples generated by the dense model. This method can accelerate the sparse training of DM and reuse the knowledge of the dense model as much as possible. This method does not change the number of sampling steps used to generate samples. The following is a formal description of progressive knowledge transfer.
${\bm{x}}_{0}$ is the data used to train DM. During training, a sparse neural network $\mathcal{F}_{\mathcal{W}}({\bm{x}}_{t},t)$ is trained to predict the noise in ${\bm{x}}_{t}$ w.r.t. ${\bm{x}}_{0}$ by minimizing the L2 loss between them. The loss is formulated as follows:
@@ 3.2 Transfer Learn Sparse Diffusion Models
The overall loss function of sparse training comprises the original dense task loss, a diffusion loss that optimizes the diffusion model. It can be formulated as:
$\mathcal{L}_{\mathrm{sparse}}^{i}=\lambda_{1}\mathcal{L}_{\mathrm{dense}}+\lambda_{2}\mathcal{L}_{\mathrm{diff}},~{}~{}i\in[1,k]$ (11)
where $\lambda_{1}$ and $\lambda_{2}$ are hyper-parameters to balance the losses, with range of $[0,1]$, and $k$ is the sparse mask number. When training a sparse model, especially extremely sparse models, we will select a teacher model, such as a dense or sparse model, and adjust the values of $\lambda_{1}$ and $\lambda_{2}$.
@@
$\mathcal{W}_{t+1}\leftarrow\mathcal{W}_{t}-\gamma_{t}g(\widetilde{\mathcal{W}^{(N_{i}:M_{i})}_{t}}),~{}~{}i\in[1,k]$ (13)
@@
$\mathcal{W}_{t+1}\leftarrow\mathcal{W}_{t}-\gamma_{t}\left(g(\widetilde{\mathcal{W}^{(N_{i}:M_{i})}_{t}})+\lambda_{W}(\mathcal{W}_{t}-\widetilde{\mathcal{W}^{(N_{i}:M_{i})}_{t}})\right),~{}~{}i\in[1,k]$ (14)
@@
The structured sparse function provides fully connected layers and convolutional layers with 2:4 sparse weights to achieve 2:4 sparse acceleration on the GPU. If their weights are pruned ahead of time, these layers can be accelerated using structured sparse functions on the GPU. To test real inference acceleration on an NVIDIA GPU, we used the 2:4 sparse operator in the acceleration library provided by NVIDIA. NVIDIA Ampere architecture GPUs have implemented CUDA operators to accelerate the multiplication of such matrices. 2:4 sparsity inference is expressed as:
@@
Transposable masks (Hubara et al., 2021) is one of the critical technologies for NVIDIA Ampere architecture GPUs to accelerate sparse matrices. Transposable masks suggest that a weight matrix and its transpose can be simply pruned by multiplying binary masks, so the backward pass shares the same sparse weight matrix with the forward pass.
@@ 4.1 Experimental Settings
Baseline Methods. We compare the performance of our method with the following baseline algorithm at 2:4 or other sparsity. Diff-Pruning (Fang et al., 2023b) is a method of structural pruning of diffusion models. ASP (Automatic SParsity) (Pool & Yu, 2021; Mishra et al., 2021) is a tool that enables 2:4 sparse training and inference for PyTorch models provided by Nvidia. Nvidia developed a simple training workflow that can easily generate a 2:4 structured sparse network that matches the accuracy of the dense network. STE-based Pruning (Bengio et al., 2013; Zhang et al., 2022) only uses sparse masks for training and inference. To compare the performance of other sparse masks, we give the results of inference with different sparsity rates. Finally, we provide the ablation study results of STE-based for the training process and traditional progressive sparse training. We give the ablation study results of untrained and STE-trained masks for the inference process.